# Bring My Cup! Personalizing Vision-Language-Action Models with Visual Attentive Prompting

### v1

@@ 4.1 Simulation Benchmarks
Existing VLA benchmarks mainly test generalization to novel categories or compositional instructions, and seldom require disambiguating a user-designated instance from visually similar distractors using reference images. To study this personalization setting in simulation, we construct Personalized-SIMPLER and Personalized-VLABench by adapting SIMPLER (li2024evaluating) and VLABench (zhang2025vlabench). In both benchmarks, we preserve the original task semantics but replace one task-relevant object with a user-specific instance, add same-category distractors, and provide a small reference set $R_{o}$ for that object. Personalized-SIMPLER focuses on manipulation across two robot embodiments, and Personalized-VLABench isolates multi-view personalized selection, (i.e., the end-effector moves to and touches the personal object). Examples are shown in Figure 2 (top).
Personalized-SIMPLER. SIMPLER provides high-fidelity manipulation environments on common robot platforms (li2024evaluating). We personalize six tasks across two single-camera settings: Fractal with the Google Robot and Bridge with the WidowX arm, where in each episode one task-relevant object is replaced by a personal object and accompanied by same-category distractors. Personal objects are instantiated with 3D assets from Sketchfab, [^1] https://sketchfab.com and for each we use five user-provided reference images to form $R_{o}$. We follow the original SIMPLER evaluation configurations for both Fractal and Bridge. For Fractal, the visual-matching and variant-aggregation tracks differ only in visual perturbations (lighting, camera pose, background), and systematically enumerating these variants yields 1,685 personalized episodes. For Bridge, the original SIMPLER configuration provides 96 evaluation episodes (Table 1).
Personalized-VLABench. VLABench is a large-scale benchmark for language-conditioned manipulation and long-horizon reasoning (zhang2025vlabench). We adapt it to five selection-style tasks on a Franka arm with three synchronized cameras. Each task defines a personal category (bags, shoes, figurines, miniatures, tableware); in each episode, one object in the scene is replaced by a user-specific instance and two same-category distractors are added. As in Personalized-SIMPLER, we use 3D models from Sketchfab and use 5 user-provided reference images of each personal object to form $R_{o}$. Episodes provide the personalized instruction (e.g., “select my leather bag”), the reference images, and the three camera views. Personalized-VLABench contains 5 tasks with 50 episodes each.
(a) The real-world hardware setup, featuring the SO-101 robot arm and a multi-view camera system.
(b) Objects used in the real-world experiments. Each of the 8 categories consists of one target personal object and two distractors.
@@ 4.2 Real-world Benchmarks
In addition to simulation, we construct a real-world personalization benchmark that follows the same structure. The physical setup is shown in Figure 4, and example tasks are visualized in Figure 2 (bottom).
Our setup consists of a tabletop workspace, a 6-DoF SO-101 robot arm [^2] https://huggingface.co/docs/lerobot/so101 with a 1-DoF parallel gripper, and three RGB cameras (front, side, and wrist). We consider eight everyday categories: vase, plushie, cup, slipper, plant, stuffed toy, marker, and dish scrubber. For each category, we collect three unseen physical instances from local stores and office environments. One is designated as the personal object and two serve as visually similar distractors.
Each episode places the three instances of a category and provides a small set of reference images of the personal object, captured from the real cameras and cropped to form $R_{o}$, together with a personalized instruction (e.g., “select my slipper” or “put my scrubber into the plastic bowl”). We include both selection and pick-and-place tasks, yielding a real-world benchmark that evaluates whether a policy adapted via VAP can reliably identify and manipulate user-specific objects on physical hardware.
Table 2: Performance on the Personalized-SIMPLER benchmark with the Google-Robot platform. VAP consistently outperforms all baselines. The two tracks evaluate matching against single objects and aggregates of visually similar objects.
@@
$\pi_{0}$ + VAP 87.3 58.2 72.4 58.2
Table 3: Performance on the Personalized-SIMPLER benchmark using the WidowX platform across four manipulation tasks. For each task, we evaluate 10 runs of 24 episodes. VAP ’s modular perception pipeline achieves consistently high success rates, whereas prior methods struggle to personalize.
@@
Table 4: Success Rates (%) on the Personalized-VLABench benchmark. For each task, we evaluate 10 runs of 25 episodes. VAP outperforms other baselines across all scenarios.
@@ 5 Experiments
We now describe the experimental setup, baselines, and metrics used to evaluate VAP on the benchmarks in Section 4. Additional experimental details, including environment configurations and hyperparameters, are provided in Appendix A.
@@ 5.1 Experimental Setup
Backbone VLA. Unless otherwise stated, we use $\pi_{0.5}$ (black2025pi05) as the backbone VLA policy. For experiments on Personalized-SIMPLER, we instead use the $\pi_{0}$ checkpoint released for the Fractal and Bridge settings in the open-source repository, [^3] https://github.com/allenzren/open-pi-zero following prior work showing that $\pi_{0}$ attains consistently strong performance on these settings (fang2025intact). We fine-tune $\pi_{0.5}$ solely for environment adaptation on generic data that excludes personal objects and personalized instructions, and then keep the backbone frozen for VAP and all baselines. In all experiments, the chosen backbone VLA is used as provided and remains frozen; VAP and all baselines share the same backbone in each setting. Further details on the training and adaptation of $\pi_{\text{VLA}}$ are given in Appendix A.2.
Evaluation Metrics. We report Success Rate (SR), the fraction of episodes that complete the task, following standard VLA evaluations (black2025pi05; kim2024openvla). For pick and pick-and-place tasks, we also report Correct Movement Ratio (CMR), the fraction of episodes that move the correct personal object at least once, regardless of final success (sridhar2025ricl). CMR is not defined for pointing or selection tasks. Higher values are better for all metrics.
Implementation. For the input-side module in VAP, we use DINOv2 (oquab2023dinov2) as the visual encoder, Grounding DINO (liu2024groundingdino) as the open-vocabulary detector, and SAM2 (ravi2024sam) for segmentation and tracking.
@@ 5.2 Baselines
We evaluate the following methods, all using the same VLA backbone in each setting.
@@
    Generic VLA: A naive baseline that uses VLA alone, feeding the personalized instruction directly to the model (e.g., “pick up my shaver”) and measuring its behavior without any personalization mechanism.
@@
    Hard Prompt (Short/Long): A language-based baseline that rewrites the personalized instruction with an LLM into a short or long generic description conditioned on a textual description of the object derived from its reference images. The resulting description is prepended to the instruction at inference time (see Appendix A.3 for details).
@@
    Soft Prompt (Token Learning): A token-learning baseline adapted from prior VLM personalization method Yo’LLaVA (nguyen2024yollavapersonalizedlanguagevision). For each personal object, we learn single token embedding in the underlying VLM from its reference images and then copy these learned embeddings into the VLA’s language embedding space. At inference time, the learned tokens are prepended to the instruction while keeping all VLA parameters fixed.
To ensure a strong Soft Prompt baseline, we train it with Yo’LLaVA-style supervision, achieving $95\%$ accuracy on a VQA-style recognition task, and by using oracle hard negatives drawn from test-time objects. We further confirm that the learned concept token transfers into the VLA with small embedding shift and yields object-centric spatial attention (Appendix B). Nevertheless, end-to-end manipulation improvements remain limited (Section 5.3, 5.4).
@@ 5.3 Results on Simulation Benchmarks
We first study how VAP affects instance-level manipulation in the simulation benchmarks. Tables 2, 3, and 4 report SR and CMR for Personalized-SIMPLER and SR for Personalized-VLABench.
On Personalized-SIMPLER (Google Robot), VAP substantially improves both correct-object interaction (CMR) and task completion (SR), with the largest gains on the pen-holder task where the generic $\pi_{0}$ and language-based baselines frequently pick up distractors. In the visual-matching track, VAP increases SR from $8.5\%$ to $60.3\%$ and CMR from $10.5\%$ to $89.2\%$, showing that the policy interacts with the correct instance much more reliably once the visual prompt is provided. Under variant aggregation, VAP still maintains similar performance (SR $58.2\%$, CMR $87.3\%$). On the bottle task, language-only hard prompts perform competitively, likely because the required manipulation is less demanding than in the pen-holder tasks, while VAP remains consistently strong.
The WidowX setting reveals an even sharper separation between VAP and baselines. In the task 3, the generic policy and hard prompts rarely manipulate the correct object, and soft prompts achieve only limited success, whereas VAP reaches CMR/SR of $82.9/71.3$. A similar pattern appears in the placement tasks: soft prompts can achieve high CMR yet fail to convert that interaction into task completion. For example, on Task 5, Soft Prompt attains CMR $79.2\%$ with SR $62.5\%$, while VAP achieves $100.0\%$ CMR and $95.0\%$ SR. The gap is larger on Task 6, where Soft Prompt reaches CMR $87.5\%$ but SR only $29.2\%$ (vs. $100.0/75.6$ with VAP). While the high CMR may partly reflect incidental object motion in a tightly cluttered setting, VAP still achieves much higher SR, showing more reliable end-to-end task completion.
[…]
Across benchmarks, Soft Prompt exhibits a recurring pattern. It can bias the policy toward the correct instance and yield moderate gains in CMR, but these gains do not reliably translate into end-to-end task completion (SR), particularly on placement tasks. We provide additional details and analysis of the token-learning baseline in Appendix B and Appendix B.3.
Figure 5: Performance on real-world tasks, averaged over 20 trials. We report Success Rate (SR) for all tasks and Correct Movement Ratio (CMR) for pick-and-place tasks. The CMR metric is not applicable for pointing tasks. VAP demonstrates a consistently high success rate across all scenarios, consistently outperforming the baseline.
@@
Init. (1st frame) Grounding DINO 0.19
Crop / embed / segment 0.07
@@
Table 5: Average runtime of the components of VAP in Personalized-SIMPLER. The initialization phase runs once at the beginning of each episode, while SAM2 tracking and VLA policy inference are executed at every control step.
@@ 5.4 Results on Real-world Benchmark
We next evaluate VAP on the real-world benchmark, which consists of four pointing tasks and four pick-and-place tasks executed with the $\pi_{0.5}$ policy. Figure 5 reports Success Rate (SR) for all tasks and Correct Movement Ratio (CMR) for pick-and-place.
@@
Personalized-SIMPLER (Fractal) Personalized-SIMPLER (Bridge) Personalized-VLABench Real-world Scenarios
@@
Table 6: Error-case of VAP across benchmarks. Fail is overall and Case 1–3 are conditional on failure and — denotes not applicable.
@@ 5.5 Error Case Analysis
While VAP consistently improves performance, it still exhibits a few recurring failure patterns across Personalized-SIMPLER, Personalized-VLABench, and our real-world benchmark. As summarized in Table 6, single-view settings are rarely limited by prompt ambiguity (Case 1 accounts for only 13.1%/21.4% of failures in Fractal/Bridge), and most failures occur even with correct prompting, indicating that manipulation difficulty dominates in these regimes. In contrast, multi-view settings are primarily affected by cross-view inconsistency (Case 2 accounts for 59.2% and 50.9% of failures in Personalized-VLABench and Real-world), where the personal object is not visible in a subset of cameras and grounding/tracking can drift—most notably under occlusion in the side view. We provide qualitative examples and further discussion of these multi-view and control-related failure modes in Appendix D.
@@ 5.6 Efficiency of VAP
A natural concern with VAP is whether the additional perception modules incur prohibitive runtime overhead compared to running the frozen VLA alone. Table 5 summarizes the average component runtimes in our Personalized-SIMPLER implementation. At the start of each episode, VAP establishes the personalized mask with a one-time initialization cost of $0.26$ s, consisting of Grounding DINO for category proposal generation ($0.19$ s) plus cropping, feature embedding, and initial segmentation ($0.07$ s). During control, the dominant cost remains VLA inference at $0.20$ s per step, while SAM2 performs mask tracking in $0.02$ s per frame, adding only a $10\%$ overhead per step. Overall, VAP adds modest wall-clock cost beyond the base policy while enabling the substantial improvements in SR and CMR reported in Sections 5.3 and 5.4.
@@ 6 Conclusion
We introduce Visual Attentive Prompting (VAP), a simple-yet-effective training-free personalization pipeline that enables frozen VLA models to manipulate user-specific objects from a few reference images by grounding the target and prompting the policy with a mask-based highlight aligned with an instruction rewrite. Across two simulation benchmarks and a real robot, VAP improves success and correct-object interaction over personalization baselines. At the same time, VAP still depends on the quality of the initial segmentation and on maintaining consistent masks across camera views, and we evaluate primarily single-object, short-horizon tasks. We therefore view our work as a core building block for richer forms of personalization, such as user-specific spatial conventions (e.g., “my side of the desk”) and personalized multi-step routines, whose robustness under more challenging real-world conditions we leave for future work.
@@ A.2 Training VLAs
Training Dataset for Real-world Scenarios. To adapt $\pi_{0.5}$ to the real-world tabletop setting, we collect a small, task-specific dataset using teleoperation on the SO-101 arm. We define five tasks: two selection tasks (“point at the cup”, “point at the cat figurine”) and three pick-and-place tasks (“put the miniature house into the plastic bowl”, “put the owl figurine into the plastic bowl”, “put the red bus into the plastic bowl”). All training objects in this dataset are disjoint from the personal objects used at evaluation time, so the policy only learns generic selection and pick-and-place skills rather than memorizing specific items.
For each task, we collect 50 teleoperated episodes at 30 Hz using the onboard RGB cameras, resulting in 250 episodes in total. As in the simulation setting, we train the policy to predict joint-space deltas between consecutive states. This dataset allows $\pi_{0.5}$ to adapt to the kinematics, camera viewpoints, and workspace of the real-world setup while keeping the personal objects unseen during training. Figure 7,  8 shows example episodes from the real-world scenario training data.
Training settings. For both Personalized-VLABench and the real-world dataset, we initialize $\pi_{0.5}$ from the public checkpoint and fine-tune it with the same hyperparameters. We use a batch size of 64, an action dimension of 32, and an action horizon of 4, and train for 60,000 gradient steps using Adam. We train on a single node with four A100 GPUs and use the final iteration as the last checkpoint.

### v2

@@ 4.1 Simulation Benchmarks
To study personalization in simulation, we construct Personalized-SIMPLER and Personalized-VLABench by adapting SIMPLER (Li et al., 2025) and VLABench (Zhang et al., 2025). In both settings, we replace an object with a user-specific instance, add same-category distractors, and provide a small reference set $R_{o}$. Personalized-SIMPLER evaluates manipulation across two robots, whereas Personalized-VLABench isolates multi-view selection (Figure 2, top).
Personalized-SIMPLER. We personalize six tasks from SIMPLER (Li et al., 2025) across the Fractal (Google Robot) and Bridge (WidowX) settings. In each episode, a task-relevant object is replaced by a Sketchfab asset (Sketchfab, 2026) accompanied by distractors, using five reference images to form $R_{o}$. Following original evaluation protocols, we generate 1,685 Fractal episodes across visual-matching and variant-aggregation tracks (enumerating visual perturbations), and use the standard 96 episodes for Bridge (Table 1).
Personalized-VLABench. We adapt VLABench (Zhang et al., 2025) to five selection tasks on a Franka arm with three synchronized cameras. Focusing on categories like bags and shoes, each episode replaces one object with a user-specific instance (using Sketchfab assets and 5 reference images) amidst two same-category distractors. The benchmark comprises 250 episodes (5 tasks $\times$ 50 episodes), providing personalized instructions (e.g., “select my leather bag”) and multi-view observations.
(a) The SO-101 tabletop setup with RGB cameras.
(b) Real-world objects: one personal instance and two distractors.
@@ 4.2 Real-world Benchmarks
We construct a real-world benchmark analogous to the simulation structure (Figure 2, bottom). The workspace features a 6-DoF SO-101 arm from the LeRobot platform (HuggingFace, 2026) and three RGB cameras (front, side, wrist). We target 8 everyday categories (e.g., cup, slipper), employing three unseen physical instances per category: one target and two lookalike distractors (Figure  4).
In each episode, we arrange the three instances and provide the system with the user’s reference set $R_{o}$—captured and cropped directly from the physical cameras—alongside a personalized instruction (e.g., “select my slipper” or put my scrubber into the bowl”). Spanning both selection and pick-and-place tasks, this benchmark rigorously evaluates whether VAP can reliably identify and manipulate user-specified objects on physical hardware.
Table 2: Performance on the Personalized-SIMPLER benchmark with the Google Robot platform. Track 1 evaluates visual matching under unseen personal objects, and Track 2 aggregates performance across systematic visual variants. VAP outperforms all baselines.
@@
$\pi_{0}$ + VAP 87.3 58.2 72.4 62.5
Table 3: Performance on the Personalized-SIMPLER benchmark using the WidowX platform across four manipulation tasks. For each task, we evaluate 10 runs of 24 episodes and report the mean. VAP’s modular perception pipeline achieves consistently high success rates, whereas prior methods struggle to personalize.
@@
Table 4: Performance on the Personalized-VLABench benchmark. For each task, we evaluate 10 runs of 50 episodes and report the mean. VAP outperforms other baselines across all scenarios.
@@ 5 Experiments
This section describes the experimental setup, metrics, and baselines used to evaluate VAP on the benchmarks in Section 4. Appendix A provides additional environment and hyperparameter details.
@@ 5.1 Experimental Setup
Backbone VLA. Unless otherwise stated, we use $\pi_{0.5}$ (Intelligence et al., 2025) as the backbone VLA policy. For experiments on Personalized-SIMPLER, we instead use the $\pi_{0}$ checkpoint released for the Fractal and Bridge settings in the open-source repository (Ren, 2026), following prior work showing that $\pi_{0}$ attains consistently strong performance on these settings (Fang et al., 2025). We fine-tune the base checkpoint $\pi_{0.5}$ solely for environment adaptation, using generic data that explicitly excludes personal objects and personalized instructions. The resulting backbone $\pi_{\text{VLA}}$ remains frozen throughout all experiments. Both VAP and the baselines share this identical policy to ensure a fair comparison. Appendix A.2 provides further details on this adaptation process.
Evaluation Metrics. We report Success Rate (SR), the fraction of episodes that complete the task, following standard VLA evaluations (Intelligence et al., 2025; Kim et al., 2024). For pick and pick-and-place tasks, we additionally report Correct Movement Ratio (CMR), the fraction of episodes in which the policy moves the target personal object at least once, regardless of final success (Sridhar et al., 2025). CMR is not defined for pointing or selection tasks, and higher is better for all metrics.
Implementation. For the perception module in VAP, we integrate state-of-the-art vision foundation models to ensure robust grounding: DINOv2 (Oquab et al., 2023) (final-layer $\ell_{2}$-normalized [CLS] token) for discriminative visual feature matching, Grounding DINO (Liu et al., 2024b) for open-vocabulary proposal, and SAM2 (Ravi et al., 2024) for consistent mask tracking across video frames.
@@ 5.2 Baselines
We compare VAP against the following baselines, all operating on the shared frozen backbone.
@@
    Generic VLA: This baseline feeds the raw personalized instruction (e.g., “pick up my shaver”) directly to the model. It serves as a lower bound, measuring the VLA’s zero-shot ability to resolve personal references without any external grounding mechanism.
@@
    Hard Prompt (Short/Long): A language-centric baseline that substitutes visual cues with text. We extract textual descriptions from the reference images and use an LLM to append these details to the instruction (generating Short or Long variants). This setup evaluates whether detailed linguistic context alone is sufficient for personalization (details in Appendix A.4).
@@
    Soft Prompt (Token Learning): A token-learning baseline adapted from Yo’LLaVA (Nguyen et al., 2024). For each personal object, we optimize a specific token embedding within the VLA’s language encoder using the reference images. During inference, we inject this learned token into the instruction sequence to represent the personal concept, while keeping the rest of the model weights frozen.
To ensure a rigorous comparison, we maximize the strength of the Soft Prompt baseline using Yo’LLaVA-style supervision and oracle hard negatives sampled from test-time objects. This optimized setup achieves $95\%$ accuracy on VQA recognition probes. We further verify that the learned token transfers to the VLA with minimal embedding shift and successfully induces object-centric attention (Appendix B). Despite these favorable conditions, improvements in end-to-end manipulation remain marginal (Sections 5.35.4).
@@ 5.3 Results on Simulation Benchmarks
We evaluate instance-level manipulation across three simulation benchmarks (Tables 2, 3, 4).
On Personalized-SIMPLER (Google Robot), VAP yields substantial gains in both correct-object interaction (CMR) and task completion (SR). The gap is most pronounced on the pen-holder task, where generic and language-based baselines frequently fail due to distractor interference. In the visual-matching track, VAP boosts SR/CMR from $8.5\%/10.5\%$ to $60.3\%/89.2\%$, demonstrating that visual prompting effectively resolves instance ambiguity. Crucially, these gains persist under variant aggregation (SR $58.2\%$, CMR $87.3\%$), confirming robustness to visual perturbations. While Hard Prompts remain competitive on the simpler bottle task, VAP delivers the most consistent performance across all scenarios.
The WidowX (Bridge) results further support the robustness of VAP on manipulation tasks. In the picking task (Task 3), baselines struggle to reliably locate the target, whereas VAP achieves $71.3\%$ SR. This advantage extends to placement tasks (Tasks 4–6), where VAP consistently outperforms all baselines, attaining $90\%$ SR on Tasks 4 and 5. While baselines occasionally register high interaction rates (CMR) in confined setups, they fail to translate this into successful task completion. Overall, VAP is the only method that consistently achieves high SR while maintaining high CMR across these manipulation tasks.
Finally, on Personalized-VLABench (Table 4), VAP demonstrates superior performance across all multi-view selection scenarios. While text-based prompts offer partial improvements by leveraging category-level semantics (e.g., distinguishing a bag from a cup), they fail to resolve instance-specific ambiguities. Consequently, VAP surpasses the strongest baseline by a significant margin (e.g., +36.8+36.8 points on the leather bag task), underscoring the necessity of explicit visual masks for robust personalization. We provide further analysis of the token-learning baseline and attention visualizations in Appendices B and B.3.
Figure 5: Real-world performance. We report SR over 20 trials for all tasks, and CMR specifically for pick-and-place tasks (inapplicable to selection). VAP achieves the highest SR/CMR across tasks. Error bars indicate the standard deviation across the four tasks.
@@
Initialization Grounding DINO 0.19
Segmentation & Embedding 0.07
@@
Table 5: Runtime analysis. The initialization runs once per episode, while tracking and policy inference execute at every control step. VAP adds minimal overhead ($0.02$ s) to the control loop.
@@ 5.4 Results on Real-world Benchmark
P-SIMPLER (Fractal) P-SIMPLER (Bridge) Personalized-VLABench Real-world Scenarios
@@
Table 6: Failure analysis across benchmarks. We report the overall failure rate for each setting. Cases 1–3 show the breakdown of specific error modes, calculated conditionally on failed episodes (— denotes not applicable).
@@ 5.5 Error Case Analysis
We categorize failure modes into three distinct types: single-view grounding errors (Case 1), multi-view inconsistency (Case 2), and control failures despite correct grounding (Case 3). Table 6 reveals distinct failure profiles depending on the setting. In single-view environments (Personalized-SIMPLER), failures are rarely caused by visual ambiguity (Case 1 $22\%$) but are instead dominated by control limitations (Case 3). This indicates that while the visual prompt accurately localizes the target, the bottleneck lies in the VLA’s physical manipulation capabilities.
Conversely, multi-view settings (Personalized-VLABench, Real-world) encounter a specific challenge: cross-view inconsistency (Case 2). Since VAP factorizes grounding and tracking per camera view for real-time inference, occlusions or partial visibility can cause per-view matching/tracking to lock onto different same-category instances across cameras. This yields contradictory mask prompts across views, which can destabilize behavior or bias actions toward the wrong object. A natural extension is to enforce explicit cross-view association/evidence fusion so that all views agree on a single target identity before prompting. We provide qualitative examples and several plausible design directions in Appendix D (see also Appendix D.2), and leave their implementation to future work.
@@ 5.6 Efficiency of VAP
We assess the computational overhead of our perception modules in Table 5. VAP introduces a negligible one-time initialization cost ($0.26$ s) to establish the personalized mask using Grounding DINO and SAM2. During online control, the computational bottleneck remains the VLA policy inference itself ($0.20$ s/step). Critically, the SAM2-based mask tracking adds only $0.02$ s per frame, incurring a minimal $10\%$ overhead to the control loop. This confirms that VAP enhances performance without compromising the real-time responsiveness of the robotic system.
@@ 6 Conclusion
In this work, we formalize the task of personalized object manipulation for VLAs, where the policy must distinguish and manipulate specific instances among visually similar distractors. To address this challenge, we propose Visual Attentive Prompting (VAP), demonstrating that an explicit visual prompting strategy can personalize frozen VLA models at inference time without any per-object parameter updates. Across diverse simulation and real-world benchmarks, our lightweight approach significantly outperforms language-based baselines in both object disambiguation and manipulation success. While promising, the current framework remains contingent on the performance of the underlying segmentation models and the consistency of masks across multi-view setups. Future work will focus on extending this paradigm to handle more complex personalization scenarios, such as user-specific spatial preferences and multi-step long-horizon tasks.
@@ A.2 Training VLAs
Training Dataset for Real-world Scenarios. To adapt $\pi_{0.5}$ to the real-world tabletop setting, we collect a small, task-specific dataset using teleoperation on the SO-101 arm. We define five tasks: two selection tasks (“point at the cup”, “point at the cat figurine”) and three pick-and-place tasks (“put the miniature house into the plastic bowl”, “put the owl figurine into the plastic bowl”, “put the toy bus into the plastic bowl”). All training objects in this dataset are disjoint from the personal objects used at evaluation time, so the policy only learns generic selection and pick-and-place skills rather than memorizing specific items.
For each task, we collect 50 teleoperated episodes at 30 Hz using the onboard RGB cameras, resulting in 250 episodes in total. As in the simulation setting, we train the policy to predict joint-space deltas between consecutive states. This dataset allows $\pi_{0.5}$ to adapt to the kinematics, camera viewpoints, and workspace of the real-world setup while keeping the personal objects unseen during training. Figure 7,  8 show example episodes from the real-world scenario training data.
A.3 Instruction Rewriting Logic
As described in Section 3.3, VAP employs instruction rewriting to align commands with visual prompts. We align our design with prior personalization frameworks, such as DreamBooth (Ruiz et al., 2023) and Yo’LLaVA (Nguyen et al., 2024), which typically rely on specific learned tokens or identifiers to explicitly trigger personalized concepts within user commands. Following this convention, we treat the specific possessive span (e.g., “my cup”) as the designated trigger for our personalization module. However, our framework is not limited to rigid templates and can be naturally extended to handle diverse natural language variations via semantic parsing (e.g., using LLMs or embedding similarity).
In this work, however, we adopt a deterministic rule-based approach to isolate the effect of visual prompting. The process proceeds as follows: We first parse the object category $c$ (e.g., “cup”) from the designated trigger “my $c$. Then, we map the visual mask’s fixed color to its text equivalent $w_{\text{color}}$ (e.g., “red”) and mechanically replace the trigger with “the $w_{\text{color}}$ $c$. This explicit mapping effectively bridges the user’s personalized intent with the visual cue provided to the VLA.
Training settings. For both Personalized-VLABench and the real-world dataset, we initialize $\pi_{0.5}$ from the public checkpoint and fine-tune it with the same hyperparameters, except for the action horizon. We use a batch size of 64, an action dimension of 32, and train for 60,000 gradient steps using Adam. We set the action horizon to 4 for Personalized-VLABench and 32 for the real-world dataset. We train on a single node with four A100 GPUs and use the final iteration as the checkpoint. Our environment-adaptation schedule (hundreds of episodes and tens of thousands of gradient steps) follows common practice in prior VLA fine-tuning on new environments.